# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

### v1

@@ Appendix B Fundamental Theorem of Multisymmetric Polynomials
With Eq. (missing) 8, a simple induction on the number of non-zero rows of $\Omega$ completes the proof: Suppose that $\Omega$ has a single non-zero rows, say $\Omega_{i}$. In this case,
@@
We now prove Eq. (missing) 8. First, consider the sum of the products obtained by taking a term for distinct row vector $X_{i}$ from each clause in the R.H.S. of Eq. (missing) 7. The sum is represented as
@@
Next, consider the remaining products, i.e., the products obtained by taking terms for the same row of $X$ from multiple clause in the L.H.S. of Eq. (missing) 7. For instance, consider the product obtained by taking $X_{i}^{\Omega_{1}}$ and $X_{i}^{\Omega_{2}}$ from the first and second clauses, respectively, and a term for distinct row vector $X_{j}\ (i\neq j)$ from each of the other clause. Then, the product is of the form of $X_{i}^{\Omega_{1}}X_{i}^{\Omega_{2}}Q$, where $Q$ is the product of remaining variables. This can be regarded as $X^{\Omega^{\prime}}$ with $\Omega^{\prime}_{i}=\Omega_{1}+\Omega_{2}$. In general, one can observe that such a product is represented as $X^{\Omega^{\prime}}$, where each non-zero row in $\Omega^{\prime}$ is the sum of a disjoint set of non-zero rows in $\Omega$ so that the number of non-zero rows in $\Omega^{\prime}$ is strictly less than the number of those in $\Omega$, and $\lvert\Omega\rvert_{1}=\lvert\Omega^{\prime}\rvert$. It is not difficult to see that the set of $\Omega^{\prime}$ induced by products obtained by expanding R.H.S. of Eq. (missing) 7 is partitioned into orbits under the action of $S_{n}$ such that $\Omega^{\prime}_{i}\to\Omega^{\prime}_{\sigma(i)}$ for each $i\in[n]$ and $\sigma\in S_{n}$. For any representative $\Omega^{\prime}$ of each orbit, the sum of the products in Eq. (missing) 7 associated with the orbit is expressed by the orbit-stabilizer theorem as
@@ Appendix C Symmetrization
where the expectation and the probability are taken over the permutation $\pi$ over $[F]$. Then, $\Pr_{\pi}{\left[x_{\pi(i_{1})j_{1}}=1\right]}=z_{j_{1}}/F$ follows from the definition of $z_{j_{1}}$. Next, we impose on $\pi$ the condition that $f\left(\pi(i_{1})\right)=j_{1},\ldots,f\left(\pi(i_{\ell-1})\right)=j_{\ell-1}$, i.e., $x_{i_{1}j_{1}}\cdots x_{i_{\ell-1}j_{\ell-1\ }}=1$. Then, there exists $F-\left(\ell-1\right)$ choices for $\pi(i_{\ell})$, since the term does not contain the pair of variable with the same first index. To satisfy $x_{\pi(i_{\ell})j_{\ell}}=1$, it must hold that $\pi(i_{\ell})\in f^{-1}\left(j_{\ell}\right)\setminus\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$. Let $s_{\ell}$ be the number of indices in $\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$ such that they are mapped via $f$ to $j_{\ell}$. Then, there are $z_{j_{\ell}}-s_{\ell}$ choices for $\pi(i_{\ell})$ satisfying $x_{\pi(i_{\ell})j_{\ell}}=1$. Therefore,

### v2

@@ Appendix B Fundamental Theorem of Multisymmetric Polynomials
With Eq. () 8, a simple induction on the number of non-zero rows of $\Omega$ completes the proof: Suppose that $\Omega$ has a single non-zero row, say $\Omega_{i}$. In this case,
@@
We now prove Eq. () 8. First, consider the sum of the products obtained by taking a term for distinct row vector $X_{i}$ from each clause in the R.H.S. of Eq. () 7. The sum is represented as
@@
Next, consider the remaining products, i.e., the products obtained by taking terms for the same row of $X$ from multiple clause in the L.H.S. of Eq. () 7. For instance, consider the product obtained by taking $X_{i}^{\Omega_{1}}$ and $X_{i}^{\Omega_{2}}$ from the first and second clauses, respectively, and a term for distinct row vector $X_{j}\ (i\neq j)$ from each of the other clause. Then, the product is of the form of $X_{i}^{\Omega_{1}}X_{i}^{\Omega_{2}}Q$, where $Q$ is the product of remaining variables. This can be regarded as $X^{\Omega^{\prime}}$ with $\Omega^{\prime}_{i}=\Omega_{1}+\Omega_{2}$. In general, one can observe that such a product is represented as $X^{\Omega^{\prime}}$, where each non-zero row in $\Omega^{\prime}$ is the sum of a disjoint set of non-zero rows in $\Omega$ so that the number of non-zero rows in $\Omega^{\prime}$ is strictly less than the number of those in $\Omega$, and $\lvert\Omega\rvert_{1}=\lvert\Omega^{\prime}\rvert$. It is not difficult to see that the set of $\Omega^{\prime}$ induced by products obtained by expanding R.H.S. of Eq. () 7 is partitioned into orbits under the action of $S_{n}$ such that $\Omega^{\prime}_{i}\to\Omega^{\prime}_{\sigma(i)}$ for each $i\in[n]$ and $\sigma\in S_{n}$. For any representative $\Omega^{\prime}$ of each orbit, the sum of the products in Eq. () 7 associated with the orbit is expressed by the orbit-stabilizer theorem as
@@ Appendix C Symmetrization
where the expectation and the probability are taken over the permutation $\pi$ over $[F]$. Then, $\Pr_{\pi}{\left[x_{\pi(i_{1})j_{1}}=1\right]}=z_{j_{1}}/F$ follows from the definition of $z_{j_{1}}$. Next, we impose on $\pi$ the condition that $f\left(\pi(i_{1})\right)=j_{1},\ldots,f\left(\pi(i_{\ell-1})\right)=j_{\ell-1}$, i.e., $x_{i_{1}j_{1}}\cdots x_{i_{\ell-1}j_{\ell-1\ }}=1$. Then, there exist $F-\left(\ell-1\right)$ choices for $\pi(i_{\ell})$, since the term does not contain the pair of variables with the same first index. To satisfy $x_{\pi(i_{\ell})j_{\ell}}=1$, it must hold that $\pi(i_{\ell})\in f^{-1}\left(j_{\ell}\right)\setminus\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$. Let $s_{\ell}$ be the number of indices in $\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$ such that they are mapped via $f$ to $j_{\ell}$. Then, there are $z_{j_{\ell}}-s_{\ell}$ choices for $\pi(i_{\ell})$ satisfying $x_{\pi(i_{\ell})j_{\ell}}=1$. Therefore,